# CompilerDream: Learning a Compiler World Model for General Code Optimization

### v2

@@ B.4 Hyperparameters
Reward smoothing $\alpha$ (Lee, Abbeel, and Lee 2024) 0.6
@@
Table 5: Hyperparameters of DreamerV3 in our experiments.
B.5 Training Model-free Baselines
We use RLlib (Liang et al. 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al. 2017) (Haj-Ali et al. 2020), A2C (Mnih et al. 2016), IMPALA (Espeholt et al. 2018), APEX (Horgan et al. 2018), and DQN (Mnih et al. 2015). All the experiments are conducted with 10 CPUs and an RTX-3090 GPU. For the PPO baseline in Section 4.4, we trained it for around 7 hours since we observed overfit to CodeContests when training for a longer time. Other model-free methods are trained for at least 10 hours. The training process uses 5 workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set. We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al. 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 6. We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set. The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 3.
@@ B.5 Training Model-free Baselines
Table 6: Hyperparameters for the PPO baseline, well-tuned on our dataset to be deviating from the default value in RLlib.
B.6 Autotuning
In Section 4.2, we extensively trained CompilerDream on the cBench dataset to evaluate its effectiveness as an autotuning method in comparison to other approaches on CompilerGym’s leaderboard. Similar to other studies, we trained CompilerDream using all programs in cBench, excluding the one called ghostscript due to its considerable size, which would significantly hinder the training process. The action space is set to the full set of 124 actions in LLVM. All other training details remained unchanged.
B.7 Guided Search
The guided search strategy we use in Section 4.2 followed the design in the PPO+Guided Search baseline. During the search, actions are sampled from the agent’s learned policy $\pi_{\psi}(a_{t}|s_{t})$ up to 45 steps, at which point a new episode begins. Code size reductions after applying each action are all recorded to calculate the maximum. To enhance exploration, we add a uniform distribution to the policy, implemented by a $5\%$ chance of selecting a uniformly random action instead of sampling from the policy. At each step, the code size reduction achieved so far is recorded for later comparison, and the elapsed time is monitored to ensure that the search is completed within the 1-minute time limit per benchmark. Compared to the guided search in PPO+Guided Search (in Table 1), we use a horizon of 45 steps instead of 200 steps. Additionally, PPO+Guided Search includes an extra search of 500 steps on the best sequence found, which we do not include, resulting in our wall time being a few seconds shorter than theirs.
B.8 Value Estimation
Here, we provide additional details about the experiment in Section 4.3. The baseline method, Coreset-NVP (Liang et al. 2023), trains a value prediction model that scores sequences from a fixed set of optimization sequences, referred to as the core set, which is obtained through an extensive search on their training set. For each program to optimize, it enumerates the first 45 passes in the top-scored sequences and validates their effectiveness through actual compilation. Specifically, sequences are processed in descending order of scores, executing each pass within a sequence. The compiler state (i.e., the IR) resets to its initial state after all passes in a sequence have been enumerated, before moving on to the next sequence. The code size reduction achieved after each pass is recorded, and the pass, along with its corresponding sequence prefix that yields the highest code size reduction, is selected. The core set in Coreset-NVP contains 50 sequences comprising a total of 625 passes, resulting in an average sequence length of 12.5 passes. Thus, roughly 3 or 4 sequences are attempted per benchmark, given the limit of 45 passes. These passes correspond to the optimization actions defined in Section 3.1, with a total of 79 different optimization passes included.
For our experiment, we set the action space to the full set of all 124 optimization passes available in LLVM when training our world model. Focusing on predicting the target value of cumulative reward after executing each optimization pass in the core set, we replace CompilerDream’s optimization agent with a policy that sequentially outputs the sequences from the core set, allowing the world model to focus on the problem dynamics driven by these sequences. This setup is also comparable to the baseline, as Coreset-NVP employs a value predictor that outputs a 50-dimension vector corresponding to the normalized value of each sequence in the core set. That prediction model is obtained through supervised training.
We train our world model using the CodeContests dataset introduced in Section 3.4, consistent with the experiments in Section 4.4. During training, we increase the weight of the reward loss to 100.0, while keeping the other hyperparameters unchanged. To align with the methodology used in Coreset-NVP, we validate the top 45 sequence prefixes of each program with the highest predicted values through actual compilation and select the best-performing sequence. The world model is trained using three different random seeds, and the average results are reported in Table 2.
B.9 Hardware and Training Time
@@
The random search baseline in Figure 3 is conducted with 4 CPUs, which is sufficient since it is a single-thread program.
@@
The random search baseline in Table 1 utilizes 80 CPUs, as specified in the write-up attached to CompilerGym’s leaderboard (Facebook 2022). Our CompilerDream + Guided Search method in Table 1 is tested on the same machine used for training CompilerDream, equipped with 64 CPUs and an RTX-3090 GPU.
We ensure that sufficient memory is available for all experiments. Memory size should have little impact on the results in all of our experiments.
B.10 Random Seeds
@@
All of our experiments that report a min-max range or standard deviation were replicated using three different random seeds. These seeds were randomly selected, and the results are generally reproducible with other seeds as well.
Appendix C Benchmarks
CompilerGym Benchmarks
In our study, we select eight benchmarks for zero-shot test and in-domain training: BLAS, cBench, CHStone, Linux, MiBench, NPB, OpenCV, and TensorFlow. These benchmarks are part of the built-in datasets provided by CompilerGym version 0.2.5. For benchmarks with a total number of programs more than 100, we use the first 50 programs as the test set, the following 50 programs as the validation set, and all of the rest as the training set. These training and validation sets are only used for in-domain training. The datasets comprising fewer than 100 programs are not applicable for in-domain training; instead, all their programs are allocated to the test set. The number of programs in each dataset after division is detailed in Table 7.
Dataset Training Split Validation Split Test Split
BLAS 200 50 50
cBench 23 N/A N/A
CHStone 12 N/A N/A
Linux 13,794 50 50
MiBench 40 N/A N/A
NPB 22 50 50
OpenCV 342 50 50
TensorFlow 1,885 50 50
Table 7: Dataset division of 8 CompilerGym benchmarks.
AI-Generated Benchmarks
@@ Appendix C Benchmarks
To further test the generalization ability of our CompilerDream agent on different programming languages, we borrow the method from FormAI (Tihanyi et al. 2023) and generate a dataset containing 50 unique Objective-C programs using GPT-3.5. We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs. We compile the generated programs using Clang without including any third-party libraries, and all programs that cannot pass compilation are discarded.
Appendix D Extended Experimental Results
@@
Table 8: Top performances of the zero-shot CompilerDream agent on individual programs. Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
D.1 Learning Curves
The zero-shot test performance of our CompilerDream agents during training is shown in Figure 7. Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
Figure 7: Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
D.2 Quantitative Results
D.3 Program Case Study
In Addition to the predicted optimization trajectory showcase in our paper, we present more example of optimization results here. Table 8 presents the agent’s top performance outcomes on various benchmark datasets. We observe that the agent indeed produces a specialized optimization strategy tailored for each program. Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -early-cse(early common subexpression elimination) and -simplifycfg, as particularly effective in code size optimization.
Figure 8: t-SNE (Van der Maaten and Hinton 2008) visualization of programs from training and test datasets.
D.4 Data Distribution Visualization
In Figure 8, we visualize the distribution of our training and test datasets. To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space. These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program. These comprehensive feature vectors are finally processed using t-SNE (Van der Maaten and Hinton 2008) for dimensionality reduction and visualization.
Figure 8 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles). This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments (Cummins et al. 2022), which shows a significant deviation from real-world applications. Nonetheless, our visualization can still not perfectly capture the transferability across datasets. For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.
@@ D.4 Data Distribution Visualization
Table 9: Quantitative results for code size reduction, corresponding to Figure 3. We report mean and standard deviation across three runs.

### v3

@@ B.3. Hyperparameters
Reward smoothing $\alpha$ (Lee et al., 2024) 0.6
@@
Model-free Baselines
We use RLlib (Liang et al., 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al., 2017) (Haj-Ali et al., 2020), A2C (Mnih et al., 2016), IMPALA (Espeholt et al., 2018), APEX (Horgan et al., 2018), and DQN (Mnih et al., 2015). We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al., 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 11. We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set. The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 4.
Table 11. Hyperparameters for the PPO baseline, well-tuned on our dataset to be deviating from the default value in RLlib.
@@
Table 12. Modified hyperparameters in our MBPO baseline.
Hyperparameters Value
MBPO target_update_interval 10
lr 1e-3
rollout_batch_size 50000
rollout_max_length 45
init_exploration_steps 3600
B.4. Hardware and Training Time
CompilerDream
@@ B.4. Hardware and Training Time
Random Search
@@
The random search baseline in Figure 4 is conducted with 4 CPUs, which is sufficient since it is a single-thread program. The random search baseline in Table 4 utilizes 80 CPUs, as specified in the write-up attached to CompilerGym’s leaderboard (Facebook, 2022). Our CompilerDream + Guided Search method in Table 4 is tested on the same machine used for training CompilerDream, equipped with 64 CPUs and an RTX-3090 GPU.
Model-free and Model-based baselines
The hardware setting is the same as used for training CompilerDream. For the PPO baseline in Section 4.4, we train for about 7 hours, as longer training leads to overfitting on CodeContests. Other model-free methods in Section 4.4 are trained for at least 10 hours. We use 5 workers for environment interaction and 4 evaluation workers to assess checkpoints on the validation set. In Section 4.5, we train MBPO for approximately 10 hours, as its evaluation score converges within this time.
B.5. Random Seeds
@@
All experiments reporting a min-max range or standard deviation are conducted with three different random seeds. The results in Table 5, Table 6, and Table 7 are also averaged over three runs with different seeds. The seeds are randomly selected, and the results are generally consistent across other random seeds as well.
B.6. Detail of Guided Search in Autotuning
Our guided search in Section 4.2 follows the PPO+Guided Search design. Actions are sampled from the learned policy for up to 45 steps per episode, recording code size reductions after each step to track the maximum. To encourage exploration, 5% of actions are sampled uniformly at random. We record the best reduction so far and monitor elapsed time to stay within the 1-minute budget per benchmark. Unlike PPO+Guided Search (Table 4), which uses a 200-step horizon and an extra 500-step search on the best sequence, we omit the latter and use only 45 steps, making our wall time slightly shorter.
B.7. Detail of Value Prediction Experiment
We provide additional details for the experiment in Section 4.3. The baseline, Coreset-NVP(Liang et al., 2023), trains a value model to score sequences from a fixed core set, obtained via extensive search on its training set. For each program, it evaluates the top-scored sequences in order, executing up to 45 passes across them. After each full sequence, the IR resets to the initial state. Code size reductions after each pass are tracked, and the best-performing pass and prefix are selected. The core set includes 50 sequences (625 passes total, 12.5 passes each on average), so around 3–4 sequences are tried per benchmark. The 79 passes used correspond to the optimization actions in Section 3.1.
In our setup, the action space includes all 124 LLVM passes. To train the world model to predict cumulative reward values for the core set passes, we apply each core set sequence to training programs and collect trajectories. We use the CodeContests dataset (Section 3.4), consistent with the experiments in Section 4.4. Reward loss weight is set to 100.0, with other hyperparameters unchanged. For evaluation, we compile the top 45 predicted prefixes per program and report the best.
B.8. AI-Generated Benchmarks
@@
To further test the generalization ability of our CompilerDream agent on different programming languages, we borrow the method from FormAI (Tihanyi et al., 2023) and generate a dataset containing 50 unique Objective-C programs using GPT-3.5. We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs. We compile the generated programs using Clang without including any third-party libraries, and all programs that cannot pass compilation are discarded.
Appendix C Extended Experimental Results
Table 13. Top performances of the zero-shot CompilerDream RL agent on individual programs.
@@ Appendix C Extended Experimental Results
C.1. Learning Curves
The performance of our CompilerDream with RL agents during training is shown in Figure 8. Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
Figure 8. Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
[…]
The learning curves for 7 out of the 8 datasets saturate quickly, while performance on the Linux dataset remains almost constant throughout the training process.
Table 14. Quantitative results for code size reduction, corresponding to Figure 4. We report mean and standard deviation across three runs.
@@ C.1. Learning Curves
C.2. Quantitative Results
C.3. Program Case Study
In Addition to the predicted optimization trajectory showcased in our paper, we present more examples of optimization results here. Table 13 presents the RL agent’s top performance outcomes on various benchmark datasets. We observe that the agent indeed produces a specialized optimization strategy tailored for each program. Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -early-cse(early common subexpression elimination), and -simplifycfg, as particularly effective in code size optimization.
Figure 9. t-SNE (Van der Maaten and Hinton, 2008) visualization of programs from training and test datasets.
[…]
In each dataset, the majority of data points are concentrated within a few clusters, with a smaller proportion dispersed across various regions. Notably, the data distribution in the CodeContest dataset overlaps with that of most other datasets.
C.4. Data Distribution Visualization
In Figure 9, we visualize the distribution of our training and test datasets. To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space. These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program. These comprehensive feature vectors are finally processed using t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction and visualization.
Figure 9 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles). This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments (Cummins et al., 2022), which shows a significant deviation from real-world applications. Nonetheless, our visualization can still not perfectly capture the transferability across datasets. For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.